A train 240m long takes 24 secs to cross a man running at 10km/hr in the same direction as the train. The speed of the train is?
Answer by lwsshak3(11628)

A train 240m long takes 24 secs to cross a man running at 10km/hr in the same direction as the train. The speed of the train is.
**
let x=speed of train
240m/24sec=240m/24sec*km/1000m*3600sec/hr
=240*3600/24*1000 km/hr
=36km/hr (difference in speed between train and runner)
speed of train=10+36=46 km/hr